Grade 11Mechanics

Motion of a particle of mass m is described by y=ut +1/2gtsq .find the force acting on particle

Sakshi
10 Years agoGrade 11

2 Answers

Piyush
10 Years ago
Just differentiate the given equation twice w.r.t t and you will get the acceleration now multiply it with mass you will get the answer.
Akansha chaurasia
7 Years ago
We know
s=ut+1/2gt^2
Now,
ds/dt=v=u+gt
also,
dv/dt=a=g
Force=mass×acceleration
Hence the required answer is 
Force=mass×g